# mO [#1]Paola

### v1

@@ 4.2 Main Results
Table 3: Model accuracies on the A-OKVQA-Ent under multi-choice evaluation strategy. $*{}^{*}$ denotes unusual results due to failure to follow multiple-choice prompts. Bold indicates the best metric under the same setting.
@@
Table 4: Model hit rates on the A-OKVQA-Ent under entity match evaluation strategy.
@@
4.3 Zero-shot Generalization Results
Although Oven-Wiki has already incorporated 14 classic datasets [^2] ImageNet21k-P [53, 51], iNaturalist2017 [60], Cars196 [28], SUN397 [66], Food101 [5], Sports100 [15], Aircraft [38], Oxford Flower [42], Google Landmarks v2 [65], VQA v2 [17], Visual7W [69], Visual Genome [29], OK-VQA [39], Text-VQA [57] from image recognition and visual question answering, along with extensive annotation efforts that ground answers to the knowledge base, we still hope to test the generalization capability of our method using additional out-of-domain datasets. As such, we manually curated a subset from the A-OKVQA [55] validation dataset. The subset does not overlap with any dataset sources in Oven-Wiki and covers all question-answer pairs in the A-OKVQA validation split whose answer can be grounded to entities in the Wikipedia knowledge base. We name the subset A-OKVQA-Ent to emphasize this distinct property. Following the design of Oven-Wiki, we divide the subset into seen and unseen splits depending on whether the answer entity is in the Oven-Wiki train set. This subset includes 478 entries out of 1145 in the A-OKVQA validation set, of which 322 are identified as seen, while the remaining 156 are classified as unseen.
To adhere to the evaluation setting of A-OKVQA, we adopted two evaluation approaches – multi-choice and entity match. In multi-choice setting, AutoVER constructed a prefix tree with four available options to guide generation, while for other models, options are numbered in the prompt and either the correct option number or an exact entity identifier match in the response is considered a true prediction. In entity match setting, a partial match of the entity text identifier within the generated response qualifies a correct prediction. To ensure equitable comparison, baselines included are other generative models comparable in size with AutoVER-7B, specifically LLaVA [37] and its improved v1.5 version [36], OpenFlamingo [3], and InstructBLIP [11]. While we expect that all models have not seen the A-OKVQA dataset so that we can assess the zero-shot generalization ability, it is important to acknowledge that LLaVA-v1.5 has been fine-tuned with 50k A-OKVQA multi-choice instructions, potentially giving it an edge.
We present results on A-OKVQA-Ent in Tab. 3 and Tab. 4, and refer readers to  Appendix 0.E for qualitative analysis on this dataset. Our method still outperforms baseline models across all evaluation settings, with the exception of multiple-choice LLaVA-v1.5 which has been tuned with in-domain instructions. Intriguingly, we find that LLaVA-v1.5 in the entity match evaluation setting still falls short of AutoVER-7B, although it has seen in-domain samples. We also observed that LLaVA-v1 and OpenFlamingo, with inadequate instruction tuning, find it challenging to comply with the multiple-choice setting, and instead generate content unrelated to the given options, leading to their significantly lower metrics. Those findings reveal that generative language models struggle to adapt well to out-of-domain instructions and that unrestricted generation is prone to hallucinations in tasks requiring precise recognition. This further emphasizes the advantages of retrieval-assisted decoding-time augmentation for such tasks.
@@ 4.4 Ablation Study and Discussion
We discuss ablation studies on AutoVER-7B covering the effectiveness of our retrieve-generate framework, constrained decoding, and hard negative mining methods in Tab. 5. Ablations are conducted with 10% of training data, as such the metrics in ablations differ from those from the main results. We name models in ablations AutoVER-7B-0.1 to indicate such discrepancy.
Retrieval Augmentation (R). We first focus on the absence of retrieval augmentation in AutoVER. Without retrieval, the model performs constrained decoding over a prefix tree composed of all Wikipedia entity identifiers instead of a set of retrieved candidates. We notice a slight performance gain on the seen subset, likely due to the gold entity not being covered in retrieval. Conversely, the integrated retrieval design is found to significantly improve the performance of AutoVER on the unseen subset from non-viable to viable. Retrieval-augmented constrained decoding narrows down the decision-making scope of the language model, allowing AutoVER to generate entity identifiers that are never trained in the model.
Table 5: Ablation study of AutoVER-7B-0.1 on Oven-Wiki Entity Split${}_{(\texttt{Val})}$. Performance of original AutoVER-7B-0.1 is reported in the blue background cell. R: whether to enable retrieval augmentation. CD: revert to greedy decoding if disabled. LoRA: parameter-efficient tuning variant of AutoVER-7B-0.1. CLS: classifier-based variant is tuned to produce the entity prediction.
R CD LoRA CLS seen unseen hm
$\checkmark$ $\checkmark$ 48.9 19.0 27.4
$\checkmark$ 50.7 0.6 1.2
46.8 0.6 1.2
$\checkmark$ 43.5 2.8 5.3
$\checkmark$ 12.8 0.1 0.2
Constrained Decoding (CD). Upon the excluding of retrieval augmentation, we proceed to disable the entire constrained decoding mechanism, reverting AutoVER to greedily decode the next token, where grounding to an external knowledge base is no longer assured. The consequent decrease in seen accuracy demonstrates that constrained decoding effectively alleviates hallucinations of the model in predicting entities. Unfortunately, since neither checkpoints of encoder-decoder variants from [23] in Table 1 and 2 nor PaLI pre-trained weights are publicly available, we are unable to assess whether our proposed decoding-time augmenting methods can bring universal improvement on generative baselines.
@@
Autoregressive Model or Classifier (CLS). One may attribute the success of AutoVER to the large model size of its underlying MLLM and accompanied expressive capacity instead of our proposed design. In response to that, we devise a [CLS] variant of AutoVER. It follows the classical design of treating a decoder-only LM as a classifier. Specifically, we introduce a new [CLS] token into the MLLM vocabulary and append this token at the end of each query image-question pair. The corresponding last-layer hidden states are then fine-tuned for classifying the query into the 20,549 existing entities in Oven-Wiki. We observed a significant performance degradation in this setting, particularly in the unseen subset. This highlights the ineffectiveness of classifier-based VQA methods in the realm of visual entity recognition, indicating that our model performance stems from more than just its size.
@@ 4.5 Case Study
We illustrate the retrieved entity candidates and model decisions on four representative image-question pairs. We note that for the same query image, questions with different intents lead to retrieval representations with semantic relevance to the intent, laying a solid foundation for the subsequent constrained generation. For instance, in the upper part of Fig. 4, the question asking for the specific model of airplane prompts the model to retrieve various kinds of airplanes. With a slight modification from “model” to “manufacturer” in the query question, the model adapts to retrieve famous airplane manufacturer brands instead of plane models. Likewise, the model is also proficient in managing queries that demand visual localization and reasoning abilities, as depicted in the lower part of Fig. 4. We refer to Appendix 0.C for the error analysis and Appendix 0.D for more case study with comparison against GPT-4V.
@@ 5 Conclusion
We present AutoVER, a compact retrieval-augmented generation framework specifically designed for visual entity recognition. Utilizing a novel constrained decoding technique, this approach effectively overcomes the challenges of low performance in recognizing out-of-domain entities while demonstrating remarkable proficiency in questions that require visually-situated reasoning. AutoVER marks a significant advancement in visual entity recognition by doubling accuracy on nearly all challenging subsets of benchmarks.
Appendix 0.A Experimental Details
Table 6: Hyperparameter settings for training of AutoVER.
Hyperparameter Value
learning rate 2e-5
weight_decay 0
batch size per device 8
effective batch size 256
learning rate strategy Cosine
optimizer AdamW
gradient clipping Disabled
We present the hyperparameter choice in Table 6. As the effectiveness of contrastive learning only reveals in a large batch size setting, we conduct all pieces of training in a batch size of 256 on 32 V100-SXM2-32GB GPUs. Utilizing DeepSpeed [50] for distributed training management, we train AutoVER-7B for 2 days and AutoVER-13B for 6.5 days with 2.5 million training samples for one epoch.
Appendix 0.B Hard Negative Groups Construction Details
We first illustrate the process of constructing kb-hard group construction. Specifically, we consider three types of relations in Wikidata for category hierarchy construction, “instance of (P31)”, “parent taxon (P171)” and “subclass of (P279)”. To obtain category hierarchy labels, we execute the SPARQL statement shown on the left of Figure 6 on Wikidata knowledge graph. To exclude generic categories, we limit the depth of traversal on the knowledge graph to 3, and manually filter generic categories333[“Wikidata metaclass”, “second-order class”, “taxon”, “classification scheme”, “metaclass”, “third-order class”, “human”, “direct anatomical metaclass”, “organisms known by a particular common name”, “Wikimedia disambiguation page”, “variable-order class”, “concept”, “food”, “artificial physical object”, “artificial object”, “physical object”, “physical substance”, “product”, “object”, “artificial physical structure”, “architectural structure”, “equipment”, “tool”, “physical location”, “Wikimedia list article”]. On the right of Figure 6, we visualize the 3-hop category hierarchy map for the entity ATR 42. Finally, we collect entities that share any parent nodes in the category hierarchy as knowledge-similar entities.
We use ViTL/16-224px [^4] https://huggingface.co/google/vit-large-patch16-224 fine-tuned on ImageNet-1k as the visual classifier in vision-hard. While we collect the entities that share the same prediction label as vision-similar entities, we filter those entities with its image top-1 confidence score lower than 0.4, those entities without an infobox image, and those prediction labels with less than 10 entities. In Figure 7, we pick some entity images that share the image prediction “Alps” and “Goose”.
Appendix 0.C Error Analysis
Figure 5: Error analysis of two query image-question pairs. Left: two query image-question pairs. Right: some of retrieved candidates. While the correct entity answer is retrieved indicated in green font, the MLLM decides to produce the wrong entity identifier depicted in red font.
We present two typical types of errors in Figure 5. Although the model is able to retrieve the correct entity, the MLLM predicts a too specific (but actually correct as it is indeed a great blue heron) entity in the upper sample. This is due to the misunderstood of intent by the MLLM, as the query does not ask for species but category. The lower sample has the MLLM make a wrong prediction on “Wetsuit” instead of “Diving suit”.
Figure 6: kb-hard group construction illustration. Left: The SPARQL query template that used for constructing the category hierarchy of entities. We only consider category labels within 3 hops to exclude general categories like “object”, “food”, etc. : Right: Visualized 3-hop category hierarchy map for the entity ATR 42. Code highlight and category visualization are from Wikidata Graph Builder.
Figure 7: Example vision-hard groups constructed with predictions from ViTL/16-224px.
Appendix 0.D More Case Study
We present 7 cases for side-by-side comparison between GPT-4V and AutoVER-7B in Figure 9. They include situations where AutoVER-7B has an advantage (first 4 rows), GPT-4V outperforms (row 5), both make accurate predictions (row 6) and both commit mistakes (row 7). We also conclude the reasons behind GPT-4V’s failures, such as its inability to ground accurately to entities, refusal to predict with insufficient clues, hallucinatory responses, struggles with disambiguating challenging entities and lack of specific domain knowledge.
Appendix 0.E Qualitative Analysis on A-OKVQA-ENT Dataset
We present a qualitative analysis in Fig. 8 for baselines reported in Sec. 4.3. For all baselines under multi-choice settings except AutoVER, we append an explicit instruction555“Please strictly follow the options you are given. Only output the option number or the option itself.” that is omitted in the figure. Still, LLaVA-v1-7B and OpenFlamingo-9B fail to follow the instructions, leading to their extremely low evaluation results. Our AutoVER does not need any prompt engineering as it is guided by a prefix tree which never allows it to generate answers out of scope. We highlight the keywords where the model exhibits strong hallucinations in the analysis.
Appendix 0.F Current Limitations
Although AutoVER demonstrates superior performance using fewer trainable parameters compared to PaLI-17B, the accuracy on the unseen subset remains suboptimal, particularly when compared with the almost identical human performance on the seen and unseen subsets. We expect future work that focuses on exploring visual entity recognition for emerging entities that never show in the training set, and even closing the gap between model and human performance. In addition, the autoregressive nature makes it challenging to predict through a single cheap vector retrieval like a bi-encoder. Instead, it requires multiple forward passes on a billion-size language model. We believe that efficient inference such as in [21] could be an interesting direction toward real-time autoregressive visual recognition systems.
Figure 8: Qualitative analysis on AOKVQA-ENT dataset. For each triplet of image, query and choices, the upper row illustrates the input prompt and response from different baselines under multi-choice evaluation strategy, while the lower row is under entity match. Background color in red indicates an incorrect answer under the specific evaluation setting and that in green denotes a correct response. Bold highlights the model is heavily hallucinating.
Figure 9: Case study on comparison between the GPT-4V response and AutoVER-7B decision. Bold indicates the entity keyword. In the GPT-4V response, red indicates a false entity keyword and green indicates a correct one. In retrieved candidates and model decisions, red indicates a false entity prediction and green denotes a correct prediction. We also include the top-$k$ position of the GPT-4V predicted entity for reference. The answer for the last query is Matteuccia.

### v2

@@ 4.2 Main Results
4.3 Zero-shot Generalization Results
Although Oven-Wiki has already incorporated 14 classic datasets [^2] ImageNet21k-P [62, 60], iNaturalist2017 [70], Cars196 [35], SUN397 [77], Food101 [6], Sports100 [19], Aircraft [44], Oxford Flower [49], Google Landmarks v2 [76], VQA v2 [21], Visual7W [81], Visual Genome [36], OK-VQA [46], Text-VQA [67] from image recognition and visual question answering, along with extensive annotation efforts that ground answers to the knowledge base, we still hope to test the generalization capability of our method using additional out-of-domain datasets. As such, we manually curated a subset from the A-OKVQA [64] validation dataset. The subset does not overlap with any dataset sources in Oven-Wiki and covers all question-answer pairs in the A-OKVQA validation split whose answer can be grounded to entities in the Wikipedia knowledge base. We name the subset A-OKVQA-Ent to emphasize this distinct property. Following the design of Oven-Wiki, we divide the subset into seen and unseen splits depending on whether the answer entity is in the Oven-Wiki train set. This subset includes 478 entries out of 1,145 in the A-OKVQA validation set, of which 322 are identified as seen, while the remaining 156 are classified as unseen.
To adhere to the evaluation setting of A-OKVQA, we adopted two evaluation approaches – multi-choice and entity match. In the multi-choice setting, AutoVER constructs a prefix tree with four available options to guide generation, while for other models, options are numbered in the prompt and either the correct option number or an exact entity identifier match in the response is considered a true prediction. In the entity match setting, a partial match of the entity text identifier within the generated response qualifies a correct prediction. To ensure equitable comparison, baselines included are other generative models comparable in size with AutoVER-7B, specifically LLaVA [43] and its improved v1.5 version [42], OpenFlamingo [4], and InstructBLIP [14]. While we expect that all models have not seen the A-OKVQA dataset so that we can assess the zero-shot generalization ability, it is important to acknowledge that LLaVA-v1.5 has been fine-tuned with 50k A-OKVQA multi-choice instructions, potentially giving it an edge. For context, we still include LLaVA-v1.5 results although they are not directly comparable.
Table 3: Model accuracies on A-OKVQA-Ent under multi-choice evaluation. Here methods are provided with multiple choices for answers in the prompt. () Denotes unusual results due to failure to adhere to the provided multiple-choice prompts. Bold indicates the best metric under the same setting.
@@ 4.3 Zero-shot Generalization Results
Table 4: Model hit rates on A-OKVQA-Ent under entity match evaluation. Here methods are not provided with multiple choices in the prompt and are deemed successful if they produce the right entity anywhere in their output.
@@
We present results on A-OKVQA-Ent in Tab. 3 and Tab. 4, and refer readers to the supplementary material for qualitative analysis on this dataset. Our method still outperforms baseline models across all evaluation settings, with the exception of multiple-choice LLaVA-v1.5 which has been tuned with in-domain instructions. Interestingly, we find that LLaVA-v1.5 in the entity match evaluation setting still falls short of AutoVER-7B, although it has seen in-domain samples. We also observed that LLaVA-v1 and OpenFlamingo, with inadequate instruction tuning, find it challenging to comply with the multiple-choice setting, and instead generate content unrelated to the given options, leading to their significantly lower metrics. Those findings reveal that generative language models struggle to adapt well to out-of-domain instructions and that unrestricted generation is prone to hallucinations in tasks requiring precise recognition. This further emphasizes the advantages of retrieval-assisted decoding-time augmentation for such tasks.
@@ 4.4 Ablation Study and Discussion
We present ablation studies exploring the effect of retrieval augmented generation, constrained decoding, and hard negative mining in Tab. 5. Ablations are conducted with 10% of the training data and as such the metrics in ablations differ from our main results and we refer to our model as AutoVER-7B-0.1.
Retrieval Augmentation. We first focus on the absence of retrieval augmentation in AutoVER. Without retrieval, the model performs constrained decoding over a prefix tree composed of all Wikipedia entity identifiers instead of a set of retrieved candidates. We notice a slight performance gain on the seen subset, likely due to the gold entity not being covered in retrieval. Conversely, the integrated retrieval design is found to significantly improve the performance of AutoVER on the unseen subset from non-viable to viable. Retrieval-augmented constrained decoding narrows down the decision-making scope of the language model, allowing AutoVER to generate entity identifiers that are never seen during training.
Table 5: Ablation study of AutoVER-7B-0.1 on Oven-Wiki entity Split${}_{(\texttt{Val})}$.
Method seen unseen hm
AutoVER-7B-0.1 48.9 19.0 27.4
+ w/o retrieval 50.7 0.6 1.2
+ w/o constrained decoding 46.8 0.6 1.2
+ w/ LoRA 43.5 2.8 5.3
AutoVER-7B-0.1-[CLS] 12.8 0.1 0.2
Constrained Decoding. Upon the excluding of retrieval augmentation, we proceed to disable the entire constrained decoding mechanism, reverting AutoVER to greedily decode the next token, where grounding to an external knowledge base is no longer assured. The consequent decrease in seen accuracy demonstrates that constrained decoding effectively alleviates hallucinations of the model in predicting entities. Unfortunately, since neither checkpoints of encoder-decoder variants from [28] in Table 1 and 2 nor PaLI pre-trained weights are publicly available, we are unable to assess whether our proposed decoding-time augmenting methods can bring universal improvement on generative baselines.
@@
Autoregressive Model or Classifier. One may attribute the success of AutoVER to the large model size of its underlying MLLM and accompanied expressive capacity instead of our proposed design. In response to that, we devise a [CLS] variant of AutoVER. It follows the classical design of treating a decoder-only LM as a classifier. Specifically, we introduce a new [CLS] token into the MLLM vocabulary and append this token at the end of each query image-question pair. The corresponding last-layer hidden states are then fine-tuned for classifying the query into the 20,549 existing entities in Oven-Wiki. We observed a significant performance degradation in this setting, particularly in the unseen subset. This highlights the ineffectiveness of classifier-based VQA methods in the realm of visual entity recognition, indicating that our model performance stems from more than just its size.
@@ 4.5 Case Study
We illustrate the retrieved entity candidates and model decisions on four representative image-question pairs. We note that for the same query image, questions with different intents lead to retrieval representations with semantic relevance to the intent, laying a solid foundation for the subsequent constrained generation. For instance, in the upper part of Fig. 4, the question asking for the specific model of airplane prompts the model to retrieve various kinds of airplanes. With a slight modification from “model” to “manufacturer” in the query question, the model adapts to retrieve famous airplane manufacturer brands instead of plane models. Likewise, the model is also proficient in managing queries that demand visual localization and reasoning abilities, as depicted in the lower part of Fig. 4. We refer to the supplementary material for the error analysis and more case studies with comparisons against GPT-4V.
@@ 5 Conclusion
We present AutoVER, a compact retrieval-augmented generation framework specifically designed for visual entity recognition. Utilizing a novel constrained decoding technique, this approach effectively overcomes the challenges of low performance in recognizing out-of-domain entities while demonstrating remarkable proficiency in questions that require visually-situated reasoning. AutoVER marks a significant advancement in visual entity recognition by doubling accuracy on nearly all challenging subsets of benchmarks. We discuss limitations in the supplementary material.
Acknowledgements. This project was partially funded by an NSF CAREER Award #2201710, and support from the Ken Kennedy Institute at Rice University. We also thank reviewers for their feedback and encouragement.